# London WC2R 2LS, UK

### v1

@@ Abstract
We demonstrate that all perturbative scale invariant heterotic sigma models with a compact target space $M^{D}$ are conformally invariant. The proof, presented in detail for up to and including two loops, utilises a geometric analogue of the $c$-theorem based on a generalisation of the Perelman’s results on geometric flows. Then, we present examples of scale invariant heterotic sigma models with target spaces that exhibit special geometry, which is characterised by the holonomy of the connection with torsion a 3-form, and explore the additional conditions that are necessary for such sigma models to be conformally invariant. For this, we find that the geometry of the target spaces is further restricted to be either conformally balanced or the a priori holonomy of the connection with torsion reduces. We identify the pattern of holonomy reduction in the cases that the holonomy is $SU(n)$ $(D=2n)$, $Sp(k)$ $D=4k)$, $G_{2}$ $(D=7)$ and $\mathrm{Spin}(7)$ $(D=8)$. We also investigate the properties of these geometries and present some examples.
@@ 1 Introduction
Heterotic sigma models apart from the metric $G$ and $B$-field couplings, which are common with the models that have been examined so far, they also have a gauge field coupling $A$ GHMR . In addition, they are chiral and so they exhibit potential sigma model manifold reparameterisation and gauge transformation chiral anomalies. However, these are cancelled with assigning an anomalous variation to the $B$ field Hullwitten . Gravitational, gauge and mixed anomalies of certain supergravity theories also cancel with a similar mechanism previously proposed by Green and Schwarz mgjs . The understanding of these anomaly cancellations is essential to preserve these symmetries in the effective theory of heterotic strings and so ensure the geometric interpretation of the theory. In turn, the invariance of the heterotic string effective action under reparameterisations of $M^{D}$, gauge transformations of $A$ and $B$-field gauge transformations is key to the generalisation of the results of PapWitten to the heterotic sigma models.
The purpose of this article is to prove a geometric analogue of the $c$-theorem and use this to explore the relation between scale and conformal invariance in the context of heterotic sigma models. In particular, we shall establish that all perturbative scale invariant heterotic sigma models with compact target space are conformally invariant. This result has already been stated in PapWitten . Here, we shall supply the required formulae for the proof up to two loops in sigma model perturbation theory and explore the conditions on the geometry of sigma model target spaces required for conformal invariance. Then, we shall explore the relationship between scale and conformal invariance for a class of heterotic sigma models that have target spaces manifolds with special geometry – for an early related work see Hull . In particular, we shall demonstrate that heterotic sigma models with target spaces $D$-dimensional manifolds, $M^{D}$, for which the holonomy of the connection [^2] Note that ^$\widehat{\nabla}_{J}V^{I}\equiv D_{J}V^{I}+\frac{1}{2}H^{I}{}_{JK}V^{K}$, where $D$ is the Levi-Civita connection of $G$ and $H$ is given in (5), i.e. ${\mathrm{d}}H\not=0$. , $\widehat{\nabla}$, with skew-symmetric torsion $H$ is included in
@@
and the curvature $F$ of $A$ satisfies an instanton-like condition are scale invariant up to two loops in perturbation theory. This choice of holonomy groups can be motivated from either spacetime or worldsheet considerations. In the former case, they arise as conditions on the geometry of $M^{D}$ for the existence of non-trivial solutions of the gravitino Killing spinor equation. The same applies for the instanton-like condition on $F$ – it arises as a condition for solving the gaugino Killing spinor equation, for a review on the geometry of supersymmetric heterotic backgrounds see GGP . In the latter case, the $\widehat{\nabla}$-covariantly constant forms [^3] Each of the holonomy groups in (1) is characterised by the existence of $\widehat{\nabla}$-covariantly constant forms on $M^{D}$, the fundamental forms of the holonomy group. on $M^{D}$ that characterise these holonomy groups generate transformations HowePap that leave invariant the heterotic sigma model action (2). The instanton-like condition on $F$ is required for the invariance of the gauge sector deLaOssa .
@@
On comparing the conditions for scale invariance with those for conformal invariance for such heterotic sigma models, we find that the requirement for conformal invariance restricts further the geometry of $M^{D}$. In particular, there must exist a $\widehat{\nabla}$-covariantly constant vector field $\widehat{V}$ on $M^{D}$ such that $\widehat{V}=\theta+2d\Phi$, where $\theta$ is an appropriate Lee form that is constructed from the metric $G$ and the $\widehat{\nabla}$-covariantly constant forms on $M^{D}$ and $\Phi$ is the dilaton. If $\widehat{V}$ vanishes, then $\theta=-2d\Phi$ and the sigma model target space $M^{D}$ is conformally balanced. The conformal balance condition is an additional geometric restriction on $M^{D}$ – it usually arises as a condition for the solution of the dilatino Killing spinor equation of heterotic string effective theory. If $\widehat{V}$ does not vanish, the holonomy of the sigma model reduces to a subgroup of the groups stated above. In the latter case, we describe the pattern of holonomy reduction and investigate some aspects of the geometry. We also give examples of such manifolds based on a principal bundle construction – these involve an adaptation and in some cases a generalisation of constructions of Kähler and hyper-Kähler manifolds with torsion in poon ; Verbitsky .
@@
In section two, we begin with a description of the invariance of the heterotic string effective action under reparameterisations of the spacetime $M^{D}$, gauge transformations of $A$ and $B$-field gauge transformations. After establishing this, we proceed with a Perelman style of argument to define the geometric $c$-function on the sigma model couplings and prove that it is decreasing along the renormalisation group flow of the heterotic sigma model. After that the proof of the statement that scale invariant heterotic sigma models with compact target spaces $M^{D}$ are conformally invariant follows as a consequence of the invariance of the geometric $c$-function under reparameterisations of $M^{D}$, gauge transformations of $A$ and $B$-field gauge transformations. In turn, this is a consequence of the invariance of the heterotic string effective action under these symmetries.
@@
In section three, we demonstrate that sigma models with target spaces, whose the holonomy of the $\widehat{\nabla}$ connection is included in the groups listed in (1), are scale invariant. The proof requires a certain choice for the spacetime connection $\widetilde{\nabla}$ that contributes in the definition of $H$ in (5). Reviewing the work of HowePap ; deLaOssa , we point out that the heterotic sigma model action (2) is invariant under the symmetries generated by the $\widehat{\nabla}$-covariantly constant forms on manifolds with special holonomy groups that include those in (1). A special case of such a symmetry is that of $(0,p)$, $p=2,4$, worldsheet supersymmetry of heterotic sigma models previously described in Hullwitten – for completeness, we also present the off-shell superfield formulation of such heterotic heterotic sigma models given in HoweGP2 .
@@
In section four, we point out that the additional requirement of conformal invariance for these models necessitates that the vector field $\widehat{V}=\theta+2{\mathrm{d}}\Phi$ to be $\widehat{\nabla}$-covariantly constant. There are two possibilities to consider. If $\widehat{V}=0$, then the target spaces are conformally balanced which is an additional restriction on $M^{D}$. The conformal balance condition together with a restriction on the holonomy of $\widehat{\nabla}$ imposes a strong restriction on the geometry of $M^{D}$. For example it is known that compact, conformally balanced Hermitian manifolds for which the holonomy of $\widehat{\nabla}$ is contained in $SU(n)$ and $H$ is closed are Calabi-Yau with $H=0$ SIGP3 . Alternatively, if $\widehat{V}\not=0$, the holonomy of $\widehat{\nabla}$ reduces further to a subgroup of the groups listed in (1). We identify the holonomy group reduction pattern and explore the local geometry of such sigma model target spaces. For this, we use the results of GLP , see also GGP for a review, where a similar holonomy reduction pattern has been observed in the context of classification of supersymmetric heterotic string backgrounds and its consequences on the geometry of a manifold have been explored. In particular, we find that the requirement of conformal invariance reduces the a priori $SU(n)$ holonomy to $\widehat{\nabla}$ to a subgroup of $SU(n-p)$ for $p=1,\dots,n-1$. Similarly, the a priori $Sp(k)$ holonomy of $\widehat{\nabla}$ reduces to $Sp(k-q)$ for $q=1,\dots,k$, while the a priori $G_{2}$ and $\mathrm{Spin}(7)$ holonomy of $\widehat{\nabla}$ reduce to a subgroup of $SU(3)$ and $G_{2}$, respectively. We also briefly explain how the $\widehat{\nabla}$-covariantly constant forms on $M^{D}$ associated with the holonomy groups (1) generate symmetries in the heterotic sigma model action (2).
@@ 2.1 Heterotic string effective action
The fields of a (0,1)-supersymmetric (heterotic) sigma-model consist of the maps $X$ from the superspace ${\mathbb{R}}^{2|1}$ with coordinates $(u,v|\vartheta^{+})$ into a Riemannian manifold $M^{D}$ with metric $G$ and the sections $\Psi$ of the vector bundle $S_{-}\otimes X^{*}E$, where $S_{-}$ is the anti-chiral spin bundle over ${\mathbb{R}}^{2|1}$ and $X^{*}E$ is the pull-back of a vector bundle over $M^{D}$ with $X$. Here $(u,v)$ are Grassmann even light-cone coordinates while $\vartheta^{+}$ is the Grassmann odd coordinate of ${\mathbb{R}}^{2|1,1}$. The action of an (0,1)-supersymmetric sigma-model with $G$, $B$ and $A$ couplings written in terms of (0,1) superfields $X$ Hullwitten is
@@
Although the classical action (2) is invariant under all these three sigma model symmetries, quantum mechanically all three are potentially anomalous moore ; Alvarez-Gaume ; Bagger . Such anomalies affect the geometric interpretation of the theory and as a consequence the invariance of the heterotic string effective theory under the above three transformations. As we have already been mentioned, the invariance of the heterotic string effective theory is essential for the Perelman style of argument for the $c$-theorem that we seek to establish. Because of this, we shall give a brief description how the effective heterotic string effective action is constructed in such a way that it is invariant under reparameterisation, frame and gauge transformations.
@@
and $\widetilde{R}$ is the curvature of the frame connection $\widetilde{\Omega}$ on $M^{D}$. The choice of $\widetilde{\nabla}$ can be arbitrary as it can be replaced with any other connection on the spacetime up to the addition of a finite local counterterm in the sigma model the effective action. In sigma models with additional symmetry, like extended supersymmetry, a particular choice for $\widetilde{\nabla}$ has to be made for consistency. We shall elaborate in section 3 on this.
@@
where $Q_{2}^{1}(\ell,\widetilde{\Omega})$ is constructed from the usual descent equations
$P_{4}(\widetilde{R})\equiv\mathrm{tr}(\widetilde{R}\wedge\widetilde{R})={\mathrm{d}}Q_{3}(\widetilde{\Omega})\Longrightarrow\delta_{\ell}P_{4}(\widetilde{R})=0={\mathrm{d}}\delta_{\ell}Q_{3}(\widetilde{\Omega})\Longrightarrow\delta_{\ell}Q_{3}={\mathrm{d}}Q_{2}^{1}(\ell,\widetilde{\Omega})~{},$ (9)
@@
$\delta_{\ell}B_{IJ}=-\frac{\alpha^{\prime}}{4}Q_{2IJ}^{1}(\ell,\widetilde{\Omega})~{},$ (10)
at first order in $\alpha^{\prime}$, i.e. in quantum theory at one loop in sigma model perturbation theory. Such an anomalous variation is part of the Green-Schwarz mechanism mgjs of cancelling reparameterisation, gauge and mixed anomalies in certain supergravity theories in ten dimensions. Noticing the sign difference in (10) and in the last term of (9), one can conclude that
@@
In the absence of other sources, the cancellation of the global anomaly also requires that $P_{4}(\widetilde{R},F)$ represents the trivial class in the cohomology of $M^{D}$ for $H$ to be globally defined. Furthermore, notice that the right hand side of the above equation vanishes provided that the vector bundle $E$ is identified with $TM^{D}$ and $F$ is chosen to be related to $\widetilde{R}$ so that the expression in (6) vanishes. This can only happen in special cases, like for example in sigma models with (1,1) worldsheet supersymmetry.
Having motivated the use of (5) for $H$ to construct the heterotic string effective action, let us turn to Euclidean signature backgrounds. Up to an overall constant that can be absorbed in an additive shift of $\Phi$, the standard action for the $G$, $B$, $A$ and $\Phi$ fields, in the order in $\alpha^{\prime}$ indicated [^4] From now on, the notation for the ${\cal O}(\alpha^{\prime 2})$ terms will be neglected. However, it will be always assumed that the formulae apply up to ${\cal O}(\alpha^{\prime 2})$ terms., is

### v2

@@ Abstract
We demonstrate that all perturbative scale invariant heterotic sigma models with a compact target space $M^{D}$ are conformally invariant. The proof, presented in detail for up to and including two loops, utilises a geometric analogue of the $c$-theorem based on a generalisation of the Perelman’s results on geometric flows. Then, we present examples of scale invariant heterotic sigma models with target spaces that exhibit special geometry, which is characterised by the holonomy of the connection with torsion a 3-form, and explore the additional conditions that are necessary for such sigma models to be conformally invariant. For this, we find that the geometry of the target spaces is further restricted to be either conformally balanced or the a priori holonomy of the connection with torsion reduces. We identify the pattern of holonomy reduction in the cases that the holonomy is $SU(n)$ $(D=2n)$, $Sp(k)$ $D=4k)$, $G_{2}$ $(D=7)$ and $\mathrm{Spin}(7)$ $(D=8)$. We also investigate the properties of these geometries and solve the conditions for conformal invariance in some cases.
@@ 1 Introduction
Heterotic sigma models, apart from the metric $G$ and $B$-field couplings that are common with the models that have been examined so far, also have a gauge field coupling $A$ GHMR . In addition, they are chiral and so they exhibit potential sigma model manifold reparameterisation and gauge transformation chiral anomalies. However, these are cancelled with assigning an anomalous variation to the $B$ field Hullwitten . Gravitational, gauge and mixed anomalies of certain supergravity theories also cancel with a similar mechanism previously proposed by Green and Schwarz mgjs . The understanding of these anomaly cancellations is essential to preserve these symmetries in the effective theory of heterotic strings and so ensure the geometric interpretation of the theory. In turn, the invariance of the heterotic string effective action under reparameterisations of $M^{D}$, gauge transformations of $A$ and $B$-field gauge transformations is key to the generalisation of the results of PapWitten to the heterotic sigma models. Aspects of geometric flows based on the beta functions of heterotic sigma models without a gauge sector, which include the set up of flow equations and the investigation of their solitons– especially in three dimensions, have been investigated by Moroianu, Murcia and Shahbazi in carlos .
The purpose of this article is to prove a geometric analogue of the $c$-theorem and use this to explore the relation between scale and conformal invariance in the context of heterotic sigma models. In particular, we shall establish that all perturbative scale invariant heterotic sigma models with compact target space are conformally invariant [^2] The analogous statement in geometry is that all steady flow solitons are gradient flow solitons.. This result has already been stated in PapWitten . Here, we shall supply the required formulae for the proof up to two loops in sigma model perturbation theory and explore the conditions on the geometry of sigma model target spaces required for conformal invariance.
Next, we shall investigate the relationship between scale and conformal invariance for a class of heterotic sigma models that have target spaces manifolds with special geometry – for an early related work see Hull . In particular, we shall demonstrate that heterotic sigma models with target spaces $D$-dimensional manifolds, $M^{D}$, for which the holonomy of the connection [^3] Note that $\widehat{\nabla}_{J}V^{I}\equiv D_{J}V^{I}+\frac{1}{2}H^{I}{}_{JK}V^{K}$, where $D$ is the Levi-Civita connection of $G$ and $H$ is given in (5), i.e. ${\mathrm{d}}H\not=0$. , $\widehat{\nabla}$, with skew-symmetric torsion $H$ is included in
@@
and the curvature $F$ of $A$ satisfies an instanton-like condition are scale invariant up to two loops in perturbation theory. Then for $M^{D}$ compact, they must also be conformally invariant – the results have been collected in table 1. This choice of holonomy groups can be motivated from either spacetime or worldsheet considerations. In the former case, they arise as conditions on the geometry of $M^{D}$ for the existence of non-trivial solutions of the gravitino Killing spinor equation. The same applies for the instanton-like conditions on $F$ – they arises as conditions for solving the gaugino Killing spinor equation (KSE), for a review on the geometry of supersymmetric heterotic backgrounds see GGP . In the latter case, the $^\widehat{\nabla}$-covariantly constant forms [^4] Each of the holonomy groups in (1) is characterised by the existence of $^\widehat{\nabla}$-covariantly constant forms on $M^{D}$. These are the fundamental forms of the holonomy group. on $M^{D}$ that characterise these holonomy groups generate transformations HowePap that leave invariant the heterotic sigma model action (2). The instanton-like conditions on $F$ are required for the invariance of the sigma model gauge sector deLaOssa under such transformations.
@@
On comparing the conditions for scale invariance with those for conformal invariance for such heterotic sigma models, we find that the requirement for conformal invariance restricts further the geometry of $M^{D}$. In particular, there must exist a $^\widehat{\nabla}$-covariantly constant vector field $\widehat{V}$ on $M^{D}$ such that $\widehat{V}=\theta+2d\Phi$, where $\theta$ is an appropriate Lee form that is constructed from the metric $G$ and the $^\widehat{\nabla}$-covariantly constant forms on $M^{D}$ and $\Phi$ is the dilaton. If $\widehat{V}$ vanishes, then $\theta=-2d\Phi$ and the sigma model target space $M^{D}$ is conformally balanced. The conformal balance condition is an additional geometric restriction on $M^{D}$ – it usually arises as a condition for the solution of the dilatino KSE of heterotic string effective theory. If $\widehat{V}$ does not vanish, the holonomy of the sigma model reduces to a subgroup of the groups stated above in (1). In the latter case, we describe the pattern of holonomy reduction and investigate some aspects of the geometry. We also solve some of the conditions for conformal invariance using a principal bundle construction – these involve an adaptation and, in some cases, a generalisation of constructions of Kähler and hyper-Kähler manifolds with torsion in poon ; Verbitsky . We also provide (accountably) many explicit examples of geometries for which the associated sigma models are conformally invariant. These are based on constructions involving group manifolds and the hyper=Kähler manifold $K_{3}$. The resulting examples include conformally balanced and non-conformally balanced manifolds as well as manifolds for which the holonomy of the connection with torsion is included in $SU(n)$, $Sp(k)$ and $G_{2}$.
@@
In section two, we begin with a description of the invariance of the heterotic string effective action under reparameterisations of the spacetime $M^{D}$, gauge transformations of $A$ and $B$-field gauge transformations. After establishing this, we proceed with a Perelman style of argument to define the geometric $c$-function on the sigma model couplings and prove that it is decreasing along the renormalisation group flow of the heterotic sigma model. After that the proof of the statement that scale invariant heterotic sigma models with compact target spaces $M^{D}$ are conformally invariant follows as a consequence of the invariance of the geometric $c$-function under reparameterisations of $M^{D}$, gauge transformations of $A$ and $B$-field gauge transformations. In turn, this invariance is a consequence of the invariance of the heterotic string effective action under these symmetries.
@@
In section three, we demonstrate that sigma models with target spaces, whose the holonomy of the $\widehat{\nabla}$ connection is included in the groups listed in (1), are scale invariant. The proof requires a certain choice for the spacetime connection $\widetilde{\nabla}$ that contributes in the definition of $H$ in (5). As a result if $M^{D}$ is compact, then these heterotic sigma models are also conformally invariant. This section includes table 1 that these results are collected. Reviewing the work of HowePap ; deLaOssa , we point out that the heterotic sigma model action (2) is invariant under the symmetries generated by the $^\widehat{\nabla}$-covariantly constant forms on manifolds with special holonomy groups that include those in (1). A special case of such a symmetry is that of $(0,p)$, $p=2,4$, worldsheet supersymmetry of heterotic sigma models previously described in Hullwitten – for completeness, we also present the off-shell superfield formulation of such heterotic heterotic sigma models given in HoweGP2 .
@@
In section four, we point out that the additional requirement of conformal invariance for these models necessitates that the vector field $\widehat{V}=\theta+2{\mathrm{d}}\Phi$ to be $^\widehat{\nabla}$-covariantly constant. There are two possibilities to consider. If $\widehat{V}=0$, then the target spaces are conformally balanced which is an additional restriction on $M^{D}$. The conformal balance condition together with a restriction on the holonomy of $\widehat{\nabla}$ impose a strong restriction on the geometry of $M^{D}$. For example it is known that compact, conformally balanced Hermitian manifolds for which the holonomy of $\widehat{\nabla}$ is contained in $SU(n)$ and $H$ is closed are Calabi-Yau with $H=0$ SIGP3 . Alternatively, if $\widehat{V}\not=0$, the holonomy of $\widehat{\nabla}$ reduces further to a subgroup of the groups listed in (1). We identify the holonomy group reduction pattern and explore the local geometry of such sigma model target spaces. For this, we use the results of GLP , see also GGP for a review, where a similar holonomy reduction pattern has been observed in the context of classification of supersymmetric heterotic string backgrounds and its consequences on the geometry of a manifold have been explored. In particular, we find that the requirement of conformal invariance reduces the a priori $SU(n)$ holonomy to $\widehat{\nabla}$ to a subgroup of SU(np)SU(n-p) for $p=1,\dots,n-1$. Similarly, the a priori $Sp(k)$ holonomy of $\widehat{\nabla}$ reduces to Sp(kq)Sp(k-q) for $q=1,\dots,k$, while the a priori $G_{2}$ and $\mathrm{Spin}(7)$ holonomy of $\widehat{\nabla}$ reduce to a subgroup of $SU(3)$ and $G_{2}$, respectively. We also briefly explain how the $^\widehat{\nabla}$-covariantly constant forms on $M^{D}$ associated with the holonomy groups (1) generate symmetries in the heterotic sigma model action (2). We conclude this section with some explicit examples. These include Wess-Zumino-Witten (WZW) models with flat but not trivial gauge sector connections whose target spaces, which are group manifolds, admit CYT and HKT structures. We also present an infinite class of conformally invariant heterotic sigma models with target space a manifold homeomorphic to $K_{3}$ and a similar example of a sigma model with target space a $G_{2}$ manifold with torsion. Moreover, we give a counterexample which illustrates that the smoothness of geometry and the compactness of the sigma model manifold are necessary and sufficient conditions for a scale invariant heterotic sigma model to be conformally invariant.
@@ 2.1 Heterotic string effective action
The fields of a (0,1)-supersymmetric (heterotic) sigma-model consist of the maps $X$ from the superspace ${\mathbb{R}}^{2|1}$ with coordinates $(u,v|\vartheta^{+})$ into a Riemannian manifold $M^{D}$ with metric $G$ and the sections $\Psi$ of the vector bundle $S_{-}\otimes X^{*}E$, where $S_{-}$ is the anti-chiral spin bundle over ${\mathbb{R}}^{2|1}$ and $X^{*}E$ is the pull-back of a vector bundle $E$ over $M^{D}$ with $X$. Here $(u,v)$ are Grassmann even light-cone coordinates while $\vartheta^{+}$ is the Grassmann odd coordinate of ${\mathbb{R}}^{2|1,1}$. The action of an (0,1)-supersymmetric sigma-model with $G$, $B$ and $A$ couplings written in terms of (0,1) superfields $X$ Hullwitten is
@@
Although the classical action (2) is invariant under all these three sigma model symmetries, quantum mechanically all three are potentially anomalous moore ; Alvarez-Gaume ; Bagger . Such anomalies affect the geometric interpretation of the theory and as a consequence the invariance of the heterotic string effective theory under the above three transformations. As we have already mentioned, the invariance of the heterotic string effective theory is essential to establish the Perelman style of argument for the $c$-theorem. Because of this, we shall give a brief description how the effective heterotic string effective action is constructed in such a way that it is invariant under reparameterisation, frame and gauge transformations.
@@
and $\widetilde{R}$ is the curvature of the frame connection $\widetilde{\Omega}$ on $M^{D}$. The choice of $\widetilde{\nabla}$ can be arbitrary as it can be replaced with any other connection on the spacetime up to the addition of a finite local counterterm in the sigma model effective action. In sigma models with additional symmetry, like extended supersymmetry, a particular choice for $\widetilde{\nabla}$ has to be made for consistency. We shall elaborate in section 3 on this.
@@
[…]
@@
[…]
at first order in $\alpha^{\prime}$, i.e. at one loop in sigma model perturbation theory. Such an anomalous variation is part of the Green-Schwarz mechanism mgjs of cancelling reparameterisation, gauge and mixed anomalies in certain supergravity theories in ten dimensions. Noticing the sign difference in (10) and in the last term of (9), one can conclude that for $H$ in (5)
@@
This equation will also be referred to as the anomalous Bianchi identity. In the absence of other sources, the cancellation of the global anomaly also requires that $P_{4}(\widetilde{R},F)$ represents the trivial class in the cohomology of $M^{D}$ for $H$ to be globally defined. Furthermore, notice that the right hand side of the above equation vanishes provided that the vector bundle $E$ is identified with $TM^{D}$ and $F$ is chosen to be related to $\widetilde{R}$ so that the expression in (6) vanishes. This can only happen in special cases, like for example in sigma models with (1,1) worldsheet supersymmetry.
Having motivated the use of (5) for $H$ to construct the heterotic string effective action, let us turn to Euclidean signature backgrounds. Up to an overall constant that can be absorbed in an additive shift of $\Phi$, the standard action for the $G$, $B$, AA and $\Phi$ fields, in the order in $\alpha^{\prime}$ indicated [^5] From now on, the notation for the ${\cal O}(\alpha^{\prime 2})$ terms will be neglected. However, it will be always assumed that the formulae apply up to ${\cal O}(\alpha^{\prime 2})$ terms., is
@@
Although we refer to $S(G,B,A,\Phi)$ as the heterotic string effective action, we keep the dimension $D$ of $M^{D}$ arbitrary instead of setting $D=10$ required by heterotic string theory. From the heterotic string theory perspective, as it stands for $D\not=10$, $S(G,B,A,\Phi)$ gives the incorrect field equation for the dilaton – the latter requires the addition of a constant term proportional to D10D-10. But as we shall see below $S(G,B,A,\Phi)$ is an intermediate step in the construction of a Perelman style argument for the heterotic sigma models. In particular, an additional constraint is required described in (17) below and the functional $\overline{S}(G,B,A)$ that is extremized differs from that of $S(G,B,A,\Phi)$. As we shall demonstrate, the new functional gives the correct renormalisation group flow equations required for the couplings of heterotic sigma models.